Two parallel plates have equal and opposite charges. When the space between the plates is evacuated, the electric field is E= 3.10×105 V/m . When the space is filled with dielectric, the electric field is E= 2.40×105 V/m .

What is the charge density on each surface of the dielectric and the dielectric constant?
Answer #1

Given,

E = 3.1 x 10^5 V/m; Ek = 2.4 x 10^5 V/m

The induced field will be:

Ei = (3.1 - 2.4) x 10^5 = 7 x 10^4 V/m

Ei = sigma/e0 => sigma = E e0

Ei = 7 x 10^4 x 8.85 x 10^-12 = 6.19 x 10^-7 C/m^2

Hence, Ei = 6.19 x 10^-7 C/m^2